use the pythagorean theorem to find the indicated quantity. Round your answer to the nearest hundredth.
14^2 = 5^2 + a^2

Answer by checkley77(12844)

14^2 = 5^2 + a^2
196=25+a^2
a^2=196-25
a^2=171
a=sqrt171
a=13.08 ans.